The speed of a car decreases from 40 m/s to 35 m/s at 12 meters. What is the kinetic energy of the car?

1
Expert's answer
2021-11-11T17:07:45-0500

Let's assume uniform decceleration. Then the acceleration of the car is given as follows (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas):


"a = \dfrac{v_f^2-v_i^2}{2d}"

where "v_f = 12m\/s, v_i = 40m\/s" are final and initial speeds, "d = 12m" is the travelled distance. The speed of the car at any given moment of time "t" is then:


"v = v_i + at = v_i + \dfrac{v_f^2-v_i^2}{2d}t"

In turn, the kinetic energy at any given moment of time is:


"K = \dfrac{mv^2}{2} = \dfrac{m}{2}\left( v_i + \dfrac{v_f^2-v_i^2}{2d}t\right)^2"

where "m" is the mass of the car. Substituting values for "m" and "t" one can obtain the numeric value of kinetic enrgy at any moment of the motion.


Answer. "K = \dfrac{m}{2}\left( v_i + \dfrac{v_f^2-v_i^2}{2d}t\right)^2".